i need help on this question and how to do step by step
y=x^2-12x-9

Answer by rfer(16322)

a=1, b=-12, c=-9
sqrt(-12^2-4*1*-9)
sqrt(144+36)
sqrt180=13.4
--------------
(+12+13.4)/2=12.7
(+12-13.4)/2=-0.7
(0,12.7)
(0.-0.7)